OpenStudy (anonymous):

Explain how you use the pattern of imaginary numbers to determine what i^18 equals

OpenStudy (earthcitizen):

1

OpenStudy (anonymous):

jus 1??? how would i explain that??

OpenStudy (earthcitizen):

\[i = \sqrt{-1}\] \[i^8 = \sqrt{(-1)^{2}(-1)^{2}(-1)^{2}} = 1\]

OpenStudy (anonymous):

Remember that i is the square root of -1. Now look at this sequence: i=sqrt(-1), i^2=-1, i^3=-sqrt(-1), i^4=1 Now from i^5 we are just multiplying 1 by i, and that makes us return to the begining of the cycle, so 18 is not a multiple of four, but 16 is, and if we look at the begining of the cycle from 16, we need to go two steps ahead, and we see that i^18=i^2=-1 @EarthCitizen thats wrong, you did for 8, its 18.

OpenStudy (earthcitizen):

yh, I thought i saw 8 my bad @ivanmlerner is correct

OpenStudy (earthcitizen):

anyways @Swageraid32 u get the gist boss ?